The length of time that it takes Ken to drive to work represents a normal distribution with a mean of 25 minutes and a standard deviation of 4.5 minutes.If Ken allows 35 minutes to get to work , what percent of the time can he except to be late?
Answer by ewatrrr(24785)

 
Hi
Population: mean is 25 and the deviation is 4.5.
P(x ≥ 35) = 1 - P(z ≤ 2.222) = 1 - .9869 = .0131 0r 1.31 % chance of being late
For the normal distribution: Below: z = 0, z = ± 1, z= ±2 , z= ±3 are plotted.
Area under the standard normal curve to the left of the particular z is P(z)
Note: z = 0 (x value: the mean) 50% of the area under the curve is to the left and 50% to the right

one standard deviation from the mean accounts for about 68% of the set
two standard deviations from the mean account for about 95%
and three standard deviations from the mean account for about 99.7%.